# The Riemann Hypothesis is false

### v34

@@
$\zeta(\sigma+it)=\sum_{n\leq x}n^{-\sigma-it}+\chi(\sigma+it)\sum_{n\leq y}n^{\sigma-1+it}+O(x^{-\sigma}\log t+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
where $0\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h$ is a constant and $|\chi(\sigma+it)|=\Big{|}\frac{\zeta(\sigma+it)}{\zeta(1-\sigma-it)}\Big{|}\asymp(t/2\pi)^{1/2-\sigma}$ [6, p.78]. Thus since $|\zeta(it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows from (4) that
@@
for $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$. For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it similarly follows from (5) by partial summation that
@@
A similar argument yields (15). Summing (9) for $\frac{T}{\log^{2}T}<n<T$ and invoking (16), (15), (14), (13) and (12) together with $\sum_{m\leq x}\Lambda(m)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430], yields
@@
Remark 4. The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need Theorem 4.11 of Titchamrsh [6] to also hold for $\sigma=\sigma_{t}\sim 0$ with an error term $\ll(\log T)^{-1}$. However, this is not possible (as can be seen from the approximate functional equation for $\zeta$). It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}\sim 0$. Indeed, he begins his argument by quoting the Euler-McLaurin summation formula for $\zeta(\sigma+it)$ (equation 3.5.3). This is initially defined for $\sigma>1$, and then is extended by holomorphic continuation to $\sigma>0$. Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh needs the assumption that $\sigma\geq\sigma_{0}>0$.

### v35

@@
$\zeta(\sigma+it)=\sum_{n\leq x}n^{-\sigma-it}+\chi(\sigma+it)\sum_{v\leq y}v^{\sigma-1+it}+O(x^{-\sigma}\log t+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
where $0\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h$ is a constant and $|\chi(\sigma+it)|=\Big{|}\frac{\zeta(\sigma+it)}{\zeta(1-\sigma-it)}\Big{|}=\Big{|}\Big{(}\frac{2\pi}{t}\Big{)}^{\sigma-\frac{1}{2}+it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{1/2-\sigma}$ [6, p.78]. Thus since $|\zeta(it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows from (4) that
@@
uniformly for $m\geq T(1+\varepsilon)$ and $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$. For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it similarly follows from (5) by partial summation that
@@
A similar argument yields (15). Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation and invoking (16), (15), (14), (13) and (12) together with $\sum_{m\leq x}\Lambda(m)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430], yields
@@
Remark 4. The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need Theorem 4.11 of Titchamrsh [6] to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log T)^{-1}$. However, this is not possible (as can be seen from the approximate functional equation for $\zeta$). It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}=o(1)$. Indeed, he begins his argument by quoting the Euler-McLaurin summation formula for $\zeta(\sigma+it)$ (equation 3.5.3). This is initially defined for $\sigma>1$, and then is extended by holomorphic continuation to $\sigma>0$. Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh assumes that $\sigma$ is bounded below by some constant $\sigma_{0}>0$.